### v1

# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

Abstract.
Driven by the ever-growing volume and decentralized nature of data, coupled with the escalating size of modern models, distributed deep learning (DDL) has been entrenched as the preferred paradigm for training. However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Remarkably, FDA achieves this without sacrificing convergence speed — in stark contrast to the trade-offs encountered in the field. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
1. Introduction
Deep Learning (DL) stands at the forefront of the most significant technological advancements in recent times, with Deep Neural Networks (DNNs) achieving remarkable success in fields such as computer vision (Kolesnikov et al., 2020), natural language processing (Devlin et al., 2019; Brown et al., 2020) and speech recognition (Amodei et al., 2015). These rapid advancements coincide with the big data era, marked by an unprecedented scale of training datasets (Sun et al., 2017; Russakovsky et al., 2015). These datasets are not only growing in size but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations, and data sovereignty laws (Kairouz et al., 2019; McMahan et al., 2016; Konečný et al., 2016). In response, Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Peteiro-Barral and Guijarro-Berdiñas, 2012; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Moritz et al., 2016; Abadi et al., 2015; Li et al., 2020).
Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local model with the global one. The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010; Dekel et al., 2010). Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Niu et al., 2011; Cui et al., 2014; Gupta et al., 2016; Mitliagkas et al., 2016; Dutta et al., 2018; Ho et al., 2013) but sometimes compromise convergence speed and model quality.
A critical challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Shi and Chu, 2017; Tang et al., 2023). This leads to a low computation-to-communication ratio (Shi et al., 2021; Lin et al., 2017). Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
The most direct method to alleviate the communication burden is to reduce the frequency of communication rounds. Local-SGD is the prime example of this approach. It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Zhang et al., 2016; Yu et al., 2018; Haddadpour et al., 2019). Although Local-SGD is effective in reducing communication while maintaining similar accuracy levels (Smith et al., 2016; Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a significant challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Zhou and Cong, 2017; Yu et al., 2018).
To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$. In (Wang and Joshi, 2018), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps. Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and prescribed an increasing sequence. These contrasting approaches underscore the multifaceted nature of communication strategies in distributed deep learning, highlighting not only the absence of a one-size-fits-all solution but also the growing need for dynamic, context-aware strategies that can continuously adapt to the specific intricacies of the learning task.
Main Contributions. Our work addresses critical efficiency challenges in DDL, particularly in communication-constraint environments, such as the ones encountered in Federated Learning (FL) (Kairouz et al., 2019) applications. We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy, massively improving communication efficiency over previous work. Our contributions are as follows:
  • FDA dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold. This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. This extensive approach—requiring over $200$K GPU hours—compensates for the stochasticity associated with training DNNs, thereby enabling us to draw more definitive conclusions.
  • We demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of reducing communication and computation, providing greatly improved trade-offs.
  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case. Furthermore, we demonstrate that FDA leads to models with excellent generalization, inherently mitigating overfitting exhibited by previous techniques.
2. Related Work
Problem formulation. Consider distributed training of deep neural networks over multiple workers (Dean et al., 2012; Li et al., 2021). In this setting, each worker has access to its own set of training data $\mathcal{D}_{k}$, and the collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss. This scenario can be effectively modeled as a distributed optimization problem, formulated as follows:
(1) $\underset{\mathbf{w}\in\mathbb{R}^{d}}{\mathrm{minimize}}\;\;F(\mathbf{w})\triangleq\frac{1}{K}\sum_{k=1}^{K}F_{k}(\mathbf{w})$
where $K$ is the number of workers and $F_{k}(\mathbf{w})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell(\mathbf{w};\zeta_{k})\right]$ is the local objective function for worker $k$. Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $w$.
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Lin et al., 2018; Hoffer et al., 2018; Shallue et al., 2018), a challenge addressed by the post-local SGD method (Lin et al., 2018), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
Federated Averaging (FedAvg) (McMahan et al., 2016) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2019). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
Accelerating convergence. An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have deployed accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2017) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2019; Reddi et al., 2021). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2019) used control-variates (in the same spirit to SVRG (Johnson and Zhang, 2013)), with significant speed-up. FedProx (Sahu et al., 2018) re-parametarized FedAvg (McMahan et al., 2016) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in the works (Ström, 2015; Aji and Heafield, 2017; Alistarh et al., 2018), and quantization techniques, which involve transmitting only quantized gradients, as detailed in studies like (Rabbat and Nowak, 2005; Shlezinger et al., 2020). Notably, these compression techniques can be effectively combined with Local-SGD methods to enhance communication-efficiency further. A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies. A comprehensive survey of compression techniques in DDL can be found in (Wang et al., 2023).
3. Federated Dynamic Averaging
Notation. At each time instance $t$, each worker $k$ independently maintains its own vector of model parameters, [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature. denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$. Let $\mathbf{\Delta w}^{(k)}_{t}$ be the update vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) encompassing the learning rate and relevant gradients. In a distributed learning step $t$, each worker $k$ applies the following update:
(2) $\mathbf{w}^{(k)}_{t+1}=\mathbf{w}^{(k)}_{t}+\mathbf{\Delta w}^{(k)}_{t}$
Let $\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ represent the average model at time $t$ and $\overline{\mathbf{w}}_{t_{0}}$ the initial model shared among the workers at the start of the current communication round. Moreover, we introduce the drift vector $\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$, that is, the change to the local model of the $k$-th worker at time $t$ since the beginning of the current round at time $t_{0}$, and $\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}=\overline{\mathbf{w}}_{t}-\overline{\mathbf{w}}_{t_{0}}$ as the average drift.
Model Variance. The model variance quantifies the dispersion or spread of the worker models around the average model. At time $t$, it is defined as:
$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\mathbf{w}_{t_{0}}+\mathbf{w}_{t_{0}}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
(3) $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=...=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
This measure provides insight into how closely aligned the workers’ models are at any given time. High variance indicates that the models are widely spread out, essentially drifting apart, leading to a lack of cohesion in the aggregated model. Conversely, a moderate-low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective. The variance plays a critical role in our approach, as it helps us gauge the state of the training process making informed decisions about round termination.
Round Invariant. We introduce the Round Invariant (RI), a condition that bounds model variance below a threshold $\Theta$:
(4) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
Our algorithm continuously monitors the model variance to maintain the RI. If the variance estimate exceeds the threshold $\Theta$ at any point $t$, the training round immediately terminates. This triggers synchronization, where all local model weights $\mathbf{w}^{(k)}_{t}$ are averaged and consolidated into the new global model $\overline{\mathbf{w}}_{t}$.
Monitoring the RI. At this point, our focus shifts to devising a method to monitor the RI, as defined in (4). We begin by defining the local state for each worker $k$, denoted by $S_{k}(t)\in\mathbb{R}^{p}$, which encapsulates information from worker $k$ at time $t$; it is updated arbitrarily. Next, we introduce the global state, $S(t)\in\mathbb{R}^{p}$, which represents the collective state of the distributed system at time $t$.
(5) $S(t)=\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)$
Lastly, we define a special non-linear function $H:\mathbb{R}^{p}\to\mathbb{R}$, with the following property:
(6) $\;H(S(t))\leq\Theta\;\Rightarrow\;\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta\quad\forall\;t$
Conceptually, if we define $S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}\times\mathbb{R}^{d}$ and $H(v,x)=v-\left\|x\right\|_{2}^{2}$, then, it immediately follows from (3), (5) that $H\left(S\left(t\right)\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ and property (6) is satisfied. Direct monitoring of the RI is hindered by the high dimensionality of $S_{k}(t)$, with $d$ potentially ranging from millions to many billions, values commonly found in DNNs. To mitigate this communication burden, we must apply dimensionality reduction to the local drifts $\mathbf{u}_{t}^{(k)}$ and identify an appropriate function $H$, accordingly. However, this reduction in information within local states causes RI monitoring to become approximate. Upcoming sections will detail two strategies–termed ”linear”, and ”sketch”–each offering a different balance between communication efficiency and approximation accuracy.
Algorithm. The proposed FDA algorithm is formalized in Algorithm 1. To streamline notation, $t$ also denotes each training round, which unfolds into three main parts:
  1. (1)
    Broadcast: Model $\mathbf{w}_{t}$ is distributed to all workers (Line 3).
  2. (2)
    Local Training: As long as we can guarantee the RI, workers continue training concurrently (Lines 4-9). Specifically, workers perform in-parallel updates to their local models $\mathbf{w}^{(k)}_{t}$ using mini-batch SGD, or some other optimization algorithm, by processing their local data (Lines 5-7). After each update, workers compute and broadcast their local states $S_{k}(t)$ (Line 8). Then, the global state $S(t)$ is calculated via (5), and using the function $H$, an estimate of the model variance is derived and compared with the threshold $\Theta$ (Line 9). Depending on this comparison, the local training stage either repeats or terminates.
  3. (3)
    Model Averaging: This stage is triggered when we can no longer guarantee the RI, i.e., $H(S(t))>\Theta$. At this point, all local models are communicated and then averaged to form $\mathbf{w}_{t+1}$, that is, the next round’s global model (Lines 10-12).
Algorithm 1 Federated Dynamic Averaging - FDA
1:Require: $S_{k}(t)$: The local state $S_{k}(t)\in\mathbb{R}\times\mathbb{R}^{p}$ with $p\ll d$
2:Require: $H(x)$: A function s.t. $H(S(t))\leq\Theta\Rightarrow\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
3:Require: $K$: The number of workers indexed by $k$
4:Require: $\Theta$: The model variance threshold
5:Require: $b$: The local mini-batch size
6:
7:Initialize $\mathbf{w}_{1}\in\mathbb{R}^{d}$
8:for each round $t=1,2,\dots$ do
9:     communicate $\mathbf{w}_{t}$ to all workers
10:     repeat
11:          for each worker $k=1,\dots,K$ in parallel do
12:               $\mathcal{B}_{k}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
13:               $\mathbf{w}^{(k)}_{t}\leftarrow\mathbf{w}^{(k)}_{t}-\eta\nabla\ell(\mathbf{w}^{(k)}_{t};\mathcal{B}_{k})$triangleright In-place
14:          communicate $S_{k}(t)$
15:     until $H(S(t))>\Theta$triangleright As per (5)
16:     for each worker $k=1,\dots,K$ in parallel do
17:          communicate $\mathbf{w}^{(k)}_{t}$
18:     $\mathbf{w}_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}^{(k)}_{t}$
19:
3.1. Linear Approximation
[…]
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
$H\left(v\,,\,x\right)=v-x^{2}$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$, since:
$H\left(S\left(t\right)\right)$ =(5)${=}H\left(\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{k=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
Furthermore, a random choice of $\xi$ is likely to perform poorly (terminate a round prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$. A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$. A heuristic choice is to take $\overline{\mathbf{u}}_{t_{0}}$, i.e., the change vector right before the current round started. All workers can estimate this without communication cost as the difference of the models at the beginning of the current and previous rounds and use $\xi=\frac{w_{t_{0}}-w_{t_{-1}}}{\left\|w_{t_{0}}-w_{t_{-1}}\right\|_{2}}.$
3.2. Sketch Approximation
A better estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005). An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix $\Xi$:
$\operatorname{sk}\left(\mathbf{v}\right)=\Xi=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
The $\operatorname{sk}(\cdot)$ operator is linear and can be computed in $\mathcal{O}\left(l\cdot d\right)$ steps. Let $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$. The function
$\mathcal{M}_{2}\left(\Xi\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$. We then define:
$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
$H\left(\mathbf{v}\,,\,\Xi\right)=\mathbf{v}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\Xi\right)$
Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least $\left(1-\delta\right)$. To prove this, we proceed as follows: we first observe $\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)\leq(1+\epsilon)\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ with probability at least $(1-\delta)$. Then:
$H\left(S\left(t\right)\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\;\text{with prob. at least}\;(1-\delta)$
4. Experiments
4.1. Setup
Platform. We employ TensorFlow (Abadi et al., 2015), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We modified TensorFlow to include our FDA variants and all competitive algorithms.
Table 1. Summary of Experiments
Hyper-Parameters Training NN d Dataset $\Theta$ b K Optimizer Algorithms LeNet-5 62K MNIST $\{0.5,1,1.5,2,3,5,7\}$ 32 { 5, 10, …, 60 } Adam FDA, Synchronous, FedAdam VGG16* 2.6M MNIST $\{20,25,30,50,75,90,100\}$ 32 { 5, 10, …, 60 } Adam FDA, Synchronous, FedAdam DenseNet121 6.9M CIFAR-10 $\{200,250,275,300,325,350,400\}$ 32 { 5, 10, …, 30 } SGD-NM FDA, Synchronous, FedAvgM DenseNet201 18M CIFAR-10 $\{350,500,600,700,800,850,900\}$ 32 { 5, 10, …, 30 } SGD-NM FDA, Synchronous, FedAvgM (fine-tuning) ConvNeXtLarge 198M CIFAR-100 $\{25,50,100,150\}$ 32 { 3, 5 } AdamW FDA, Synchronous
Datasets & Models. The core experiments involve training Convolutional Neural Networks (CNNs) of varying sizes and complexities on two datasets: MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2012). For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2014), denoted as VGG16*, consisting of 2.6 million parameters. VGG16* was specifically adapted for the MNIST dataset, a less demanding learning problem compared to ImageNet (Russakovsky et al., 2015), for which VGG16 was designed. In VGG16*, we omitted the $512$-channel convolutional blocks and downscaled the final two fully connected (FC) layers from $4096$ to $512$ units each. Both models use Glorot uniform initialization (Glorot and Bengio, 2010). For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2016), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2016). The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021; Ridnik et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2020). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016; Zhuang et al., 2019), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Yosinski et al., 2014).
Algorithms. We consider five distributed deep learning algorithms: LinearFDA, SketchFDA, Synchronous  [^2] The name was derived from the Bulk Synchronous Parallel approach; can be understood as a special case of the FDA Algorithm 1 where $\Theta$ is set to zero., FedAdam (Reddi et al., 2021), and FedAvgM (Hsu et al., 2019); the first three are standard in all experiments. Depending on the local optimizer, Adam (Kingma and Ba, 2017) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their federated counterparts FedAdam or FedAvgM, respectively.
Evaluation Methodology. Comparing DDL algorithms is not straightforward. For example, comparing DDL algorithms based on the average cost of a training epoch can be misleading, as it does not consider the effects on the trained model’s quality. To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in the Figures). Based on this definition, we focus on two performance metrics:
  1. (1)
    Communication cost, which is the total data (in bytes) transmitted by all workers. Translating this cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers. Its impact is larger in FL scenarios, where workers often use slower Wi-Fi connections.
  2. (2)
    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in the Figures) performed by each worker. Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the nodes. Its impact is larger for nodes with lower computational resources.
Hyper-Parameters & Optimizers. Hyper-parameters unique to each training dataset and model are detailed in Table 1; $\Theta$ is pertinent to FDA algorithms and not applicable to others. Notably, a guideline for setting the parameter $\Theta$ is provided in Section 4.2. For experiments involving FedAvgM and FedAdam, we use $E=1$ local epochs, following (Reddi et al., 2021). For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2017). In these cases, FedAdam also adheres to the default settings for both local and server optimization (Reddi et al., 2021; Chollet et al., 2015). For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2016). For FedAvgM, local optimization is conducted with default settings (Hsu et al., 2019; Chollet et al., 2015), while server optimization employs SGD with momentum, setting the momentum parameter and learning rate to $0.9$ and $0.316$, respectively (Reddi et al., 2021). Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2017), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
Data Distribution. In all experiments, the training dataset is divided into approximately equal parts among the workers. To assess the impact of data heterogeneity, we explore three scenarios:
  1. (1)
    IID — Independent and identically distributed.
  2. (2)
    Non-IID: $X$— A portion $X\%$ of the dataset is sorted by label and sequentially allocated to workers, with the remainder distributed in an IID fashion.
  3. (3)
    Non-IID: Label $Y$ — All samples from label $Y$ are assigned to a few workers, while the rest are distributed in an IID manner.
Main Findings. The main findings of our experimental analyses are: (1) LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance. (2) LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation. (3) The performance of LinearFDA and SketchFDA is comparable in most experiments. SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to a number). SketchFDA significantly outperforms LinearFDA at the transfer learning scenario. (4) The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
4.2. Results
Due to the extensive set of unique experiments (over $1000$), as detailed in Table 1, we leverage Kernel Density Estimation (KDE) plots (Waskom, 2021) to visualize the bivariate distribution of computation and communication costs incurred by each strategy for attaining the Accuracy Target. These KDE plots provide a high-level overview of the cost trade-off for training accurate models. As an illustrative example, Figure 1 depicts the strategies’ bivariate distribution for the LeNet-5 model trained on MNIST with different data heterogeneity setups. In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 1) that attained the Accuracy Target of $0.985$.
Figure 1. LeNet-5 on MNIST
Figure 2. VGG16* on MNIST
Figure 3. DenseNet121 on CIFAR-10
Figure 4. DenseNet201 on CIFAR-10
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 12,  3, and 4 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 1-4, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 1-4. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 2 for all three data heterogeneity settings. For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  34), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
FDA is resilient to data heterogeneity. In DDL, data heterogeneity is a prevalent challenge, reflecting the complexity of real-world applications where the IID assumption often does not hold. The ability of DDL algorithms to maintain consistent performance in the face of non-IID data is a critical metric for their effectiveness and adaptability. Our empirical investigation reveals the FDA methods’ noteworthy resilience in such heterogeneous environments. For LeNet-5 on MNIST, as illustrated in Figure 1, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label ”00”, $60\%$). Similarly, for VGG16* on MNIST, Figure 2 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label ”00”, Label ”$8$”); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
FDA generalizes better. The factors determining how well a DL algorithm performs are its ability to: (1) make the training accuracy high, and (2) make the gap between training and test accuracy small. These two factors correspond to the two central challenges in DL: underfitting and overfitting (Goodfellow et al., 2016). For DenseNet121 on CIFAR-10, with a test accuracy target of $0.8$, as illustrated in Figure 5, Synchronous and FedAvgM exhibit overfitting, with a noticeable discrepancy between training and test accuracy. In stark contrast, the FDA methods have an almost zero accuracy gap. Please note that LinearFDA and SketchFDA reach the test accuracy target of $0.8$ much earlier (at epochs 86 and 91, respectively). Turning our focus to DenseNet201 on CIFAR-10, with a test accuracy target of $0.78$, Synchronous again tends towards overfitting, while FedAvgM shows a slight improvement but still does not match the FDA methods, which continue to exhibit exceptional generalization capabilities, evidenced by a minimal training-test accuracy gap, as shown at Figure 5. Notably, given the necessity to fix hyper-parameters $\Theta$ and $K$ for the training accuracy plots, we selected two representative examples. The patterns of performance we highlighted are consistent across most of the conducted tests.
Figure 5. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup. This unpredictability is starkly illustrated with VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Specifically, Figure 6 demonstrates (at its upper part) that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication (shown at the upper part of the figure). These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (upper part of Figures 78), the expected behavior starts to emerge. Specifically, for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers.
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 67 and 8). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 7 and 8. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
Figure 6. VGG16* on MNIST — Accuracy Target: $0.994$
Figure 7. DenseNet121 on CIFAR-10 — Accuracy Target: $0.8$
Figure 8. DenseNet201 on CIFAR-10 — Accuracy Target: $0.78$
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 67, and 8. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 1), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
To assist researchers in selecting an optimal variance threshold, Figure 9 presents empirical estimations for $\Theta$ across three distinct learning settings: FL (assuming a common channel of $0.5$Gbps), Balanced (communication-computation equilibrium), and HPC.
$\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
$\Theta_{B}$ $=3.89\cdot 10^{-5}\cdot d$
$\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
Figure 9. Empirical Estimation of the Variance Threshold
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 123, and  4. This suggests that the precision of the variance approximation is not critical; occasional ”unnecessary” synchronizations do not significantly impact overall performance. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 10).
In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, particularly if communication-efficiency is paramount.
Figure 10. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
5. Conclusions
In this paper, we introduced Federated Dynamic Averaging (FDA), an innovative, adaptive and communication-efficient algorithm for distributed deep learning. Essentially, FDA makes informed, dynamic decisions on when to terminate training rounds based on approximations of the model variance. Through extensive experiments across diverse datasets and learning tasks, we demonstrated that FDA significantly reduces communication overhead (often by orders of magnitude) without a corresponding increase in computation or compromise in model performance — contrary to the typical trade-offs encountered in the literature. Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates overfitting. Our results push the limits of modern communication-efficient distributed deep learning, paving the way for scalable, dynamic, and broadly applicable strategies.

### v2

@@
 Abstract.
-Driven by the ever-growing volume and decentralized nature of data, coupled with the escalating size of modern models, distributed deep learning (DDL) has been entrenched as the preferred paradigm for training.
+Driven by the ever-growing volume and decentralized nature of data, coupled with the need to harness this data and generate knowledge from it, has led to the extensive use of distributed deep learning (DDL) techniques for training.
+These techniques rely on local training that is performed at the distributed nodes based on locally collected data, followed by a periodic synchronization process that combines these models to create a global model.
 However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability.
 Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules.
+These drawbacks also have a direct impact on the time required for the training process, necessitating excessive time for data communication.
 To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance.
+In essence, the costly synchronization step is triggered only if the local models, which are initialized from a common global model after each synchronization, have significantly diverged.
+This decision is facilitated by the communication of a small local state from each distributed node/worker.
 Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms.
-Remarkably, FDA achieves this without sacrificing convergence speed — in stark contrast to the trade-offs encountered in the field.
 Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
+Federated Learning, Distributed Learning
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 26th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Parallel computing methodologies
 1. Introduction
-Deep Learning (DL) stands at the forefront of the most significant technological advancements in recent times, with Deep Neural Networks (DNNs) achieving remarkable success in fields such as computer vision (Kolesnikov et al., 2020), natural language processing (Devlin et al., 2019; Brown et al., 2020) and speech recognition (Amodei et al., 2015).
-These rapid advancements coincide with the big data era, marked by an unprecedented scale of training datasets (Sun et al., 2017; Russakovsky et al., 2015).
-These datasets are not only growing in size but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations, and data sovereignty laws (Kairouz et al., 2019; McMahan et al., 2016; Konečný et al., 2016).
-In response, Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Peteiro-Barral and Guijarro-Berdiñas, 2012; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Moritz et al., 2016; Abadi et al., 2015; Li et al., 2020).
-Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local model with the global one.
-The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010; Dekel et al., 2010).
-Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Niu et al., 2011; Cui et al., 2014; Gupta et al., 2016; Mitliagkas et al., 2016; Dutta et al., 2018; Ho et al., 2013) but sometimes compromise convergence speed and model quality.
-A critical challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Shi and Chu, 2017; Tang et al., 2023).
-This leads to a low computation-to-communication ratio (Shi et al., 2021; Lin et al., 2017).
+The big data era has been marked by an unprecedented scale of training datasets (Zhang et al., 2021; Rauf et al., 2024).
+These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations (especially in federated settings, such as drones collecting and collaboratively building a global model/view of an area), and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024).
+Such constraints complicate the use of Deep Learning (DL) techniques in the aforementioned scenarios, and provide challenges in harnessing knowledge from distributed data.
+Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024).
+Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local models with the global one.
+The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010).
+Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Miao et al., 2021; Fu et al., 2022) but compromise convergence speed and model quality.
+A significant challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Tang et al., 2023; Wang et al., 2016).
+This leads to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022).
 Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
@@
 Local-SGD is the prime example of this approach.
-It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Zhang et al., 2016; Yu et al., 2018; Haddadpour et al., 2019).
-Although Local-SGD is effective in reducing communication while maintaining similar accuracy levels (Smith et al., 2016; Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a significant challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Zhou and Cong, 2017; Yu et al., 2018).
+It allows workers to perform $\tau$ local update steps on their models before aggregating them, as opposed to averaging the updates in every step (Yu et al., 2019; Haddadpour et al., 2019).
+Although Local-SGD is effective in reducing communication while maintaining comparable model quality (Wang and Joshi, 2021), determining the optimal value of $\tau$ presents a critical challenge, with only a handful of studies offering theoretical insights into its influence on convergence (Wang and Joshi, 2021; Stich, 2019; Yu et al., 2019).
 To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$.
-In (Wang and Joshi, 2018), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps.
-Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and prescribed an increasing sequence.
+In (Wang and Joshi, [n.
+d.]), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps.
+Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and an increasing sequence emerged.
 These contrasting approaches underscore the multifaceted nature of communication strategies in distributed deep learning, highlighting not only the absence of a one-size-fits-all solution but also the growing need for dynamic, context-aware strategies that can continuously adapt to the specific intricacies of the learning task.
-Main Contributions.
-Our work addresses critical efficiency challenges in DDL, particularly in communication-constraint environments, such as the ones encountered in Federated Learning (FL) (Kairouz et al., 2019) applications.
-We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy, massively improving communication efficiency over previous work.
-Our contributions are as follows:
-  • FDA dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold.
+Main Idea and Contributions.
+Our work addresses critical efficiency challenges in DDL, particularly in communication-constrained environments, such as the ones encountered in Federated Learning (FL) applications (Kairouz et al., 2021).
+We introduce Federated Dynamic Averaging (FDA), a novel, adaptive distributed deep learning strategy that massively improves communication efficiency over previous work.
+FDA utilizes a novel 2-action, conditional synchronization protocol, designed to avoid the need to decide or guess the proper values of local update steps, or to synchronize after each training step, but rather only performs the costly synchronization process when needed.
+Our FDA algorithm dynamically triggers synchronization based on the value of model variance across nodes.
+In a nutshell, the costly synchronization step is only triggered if the local models have diverged significantly, which implies that the global model may no longer be accurate.
+As Figure 1 demonstrates, at the start, nodes enter the local training step with the same global model (Figure 1.(A)).
+Then, local training commences and each distributed node/worker computes its local state (Figure 1.(B)).
+This is followed by the transmission (Figure 1.(C)) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size.
+During transmission, the local states are aggregated and their average is made available to all nodes—an operation known as AllReduce.
+This operation does not require (or prohibit) the use of a central node.
+Based on the aggregated state, the nodes can estimate (Figure 1.(D)) whether the variance of the local models may have exceeded a threshold.
+If this is not the case, the costly synchronization step (Figure 1.(E)) is avoided and local training continues.
+What is important is how to properly pick these local states computed at, and then transmitted by, the local nodes.
+To address this problem, we propose two variants of our FDA algorithm.
+Our contributions can be summarized as follows:
+  • We propose FDA, an algorithm that dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold.
 This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
+  • We propose two variants of FDA, which differ on the local state that is computed at, and then transmitted by, each distributed node/worker. These two variants are termed SketchFDA and LinearFDA.
   • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks.
-This extensive approach—requiring over $200$K GPU hours—compensates for the stochasticity associated with training DNNs, thereby enabling us to draw more definitive conclusions.
-  • We demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance.
-Furthermore, it effectively balances the competing demands of reducing communication and computation, providing greatly improved trade-offs.
+Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance.
+Furthermore, it effectively balances the competing demands of communication and computation, providing greatly improved trade-offs.
   • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case.
-Furthermore, we demonstrate that FDA leads to models with excellent generalization, inherently mitigating overfitting exhibited by previous techniques.
+We note that in Federated Learning scenarios, the norm is that data are not independently and identically distributed across nodes.
+Outline.
+The remainder of this paper is organized as follows: Section 2 reviews related work.
+Section 3 introduces our DDL technique, Federated Dynamic Averaging (FDA), and its two variants.
+Section 4 details the experimental setup, and discusses the insights and conclusions drawn from our empirical investigation.
+Lastly, Section 5 contains concluding remarks.
+Figure 1.
+FDA.
+The local training step is followed by the computation of a local state by all nodes.
+Then, the (small in size) local states are aggregated.
+Based on the aggregated result, all nodes estimate if synchronization is required.
+In most cases, the expensive synchronization step of the models is avoided and local training continues
 2. Related Work
@@
 where $K$ is the number of workers and $F_{k}(\mathbf{w})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell(\mathbf{w};\zeta_{k})\right]$ is the local objective function for worker $k$.
-Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $w$.
+Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $\mathbf{w}$.
 Communication efficient Local-SGD.
@@
 The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD.
-However, the large-batch approach leads to poor generalization (Lin et al., 2018; Hoffer et al., 2018; Shallue et al., 2018), a challenge addressed by the post-local SGD method (Lin et al., 2018), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps.
+However, the large-batch approach leads to poor generalization (Hoffer et al., 2017), a challenge addressed by the post-local SGD method (Lin et al., 2020), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps.
 In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
-Federated Averaging (FedAvg) (McMahan et al., 2016) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2019).
+Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021).
 In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints.
@@
 Consequently, recent works have deployed accelerated versions of SGD to the distributed setting.
-Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2017) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a).
+Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2015) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a).
 Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting.
 However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings.
-When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2019; Reddi et al., 2021).
+When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2020; Reddi et al., 2021; Wenig and Papenbrock, 2022).
 The result is slow and unstable convergence with significant communication overhead.
-To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2019) used control-variates (in the same spirit to SVRG (Johnson and Zhang, 2013)), with significant speed-up.
-FedProx (Sahu et al., 2018) re-parametarized FedAvg (McMahan et al., 2016) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model.
+To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2020) used control-variates (in the same spirit to SVRG), with significant speed-up.
+FedProx (Sahu et al., 2018) re-parameterized FedAvg (McMahan et al., 2017) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model.
 Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
@@
 To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes.
-Key strategies include sparsification, where only crucial components of information are transmitted, as explored in the works (Ström, 2015; Aji and Heafield, 2017; Alistarh et al., 2018), and quantization techniques, which involve transmitting only quantized gradients, as detailed in studies like (Rabbat and Nowak, 2005; Shlezinger et al., 2020).
-Notably, these compression techniques can be effectively combined with Local-SGD methods to enhance communication-efficiency further.
+Key strategies include sparsification, where only crucial components of information are transmitted, as explored in (Aji and Heafield, 2017), and quantization techniques, which involve transmitting only quantized gradients, as detailed in (Shlezinger et al., 2021).
+Notably, these compression techniques can be combined with Local-SGD methods to enhance communication-efficiency further.
 A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies.
+Additionally, sketching emerges as another fundamental tool in large-scale machine learning.
+It effectively compresses high-dimensional problems into lower dimensions to save runtime and memory, typically utilizing hash-based probabilistic data structures.
+For instance, (Spring et al., 2019) use Count Sketches to compress auxiliary variables in optimization algorithms, significantly freeing up memory.
+Similarly, FetchSGD (Rothchild et al., 2020) employs Count Sketches to compress model updates and leverages their linearity for efficient merging.
+In contrast to these applications, our approach utilizes sketches not for compression but to estimate local state information, and based on this to decide whether a synchronization is required—an orthogonal application to traditional use cases.
 A comprehensive survey of compression techniques in DDL can be found in (Wang et al., 2023).
 3. Federated Dynamic Averaging
+We now present our algorithms, based on our notion of Federated Dynamic Averaging (FDA). Our algorithms deviate from prior work in these two key ways:
+  1. (1)
+    The decision on when to synchronize.
+  2. (2)
+    The actual synchronization process.
+To the best of our knowledge, this is the first Distributed Deep Learning algorithm that dynamically decides when to synchronize based on the current collective state of the training progress—whether it is advancing well or poorly.
 Notation.
-At each time instance $t$, each worker $k$ independently maintains its own vector of model parameters, [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature.
-denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$.
-Let $\mathbf{\Delta w}^{(k)}_{t}$ be the update vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) encompassing the learning rate and relevant gradients.
-In a distributed learning step $t$, each worker $k$ applies the following update:
-(2) $\mathbf{w}^{(k)}_{t+1}=\mathbf{w}^{(k)}_{t}+\mathbf{\Delta w}^{(k)}_{t}$
-Let $\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ represent the average model at time $t$ and $\overline{\mathbf{w}}_{t_{0}}$ the initial model shared among the workers at the start of the current communication round.
-Moreover, we introduce the drift vector $\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$, that is, the change to the local model of the $k$-th worker at time $t$ since the beginning of the current round at time $t_{0}$, and $\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}=\overline{\mathbf{w}}_{t}-\overline{\mathbf{w}}_{t_{0}}$ as the average drift.
-Model Variance. The model variance quantifies the dispersion or spread of the worker models around the average model. At time $t$, it is defined as:
-$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\mathbf{w}_{t_{0}}+\mathbf{w}_{t_{0}}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
-(3) $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=...=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+At each time step $t$, each worker $k$ independently maintains its own vector of model parameters [^1] The terms “model” and ”model parameters” are used interchangeably, as is common in the literature., denoted as $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$.
+Let $\mathbf{w}_{t}$ represent the $K\times d$ tensor of all local model vectors, and $\overline{\mathbf{w}}_{t}$ be the average model vector (this notation applies to all vector quantities):
+$\mathbf{w}_{t}=\left[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}\right]\quad,\quad\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$
+Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data.
+It incorporates the learning rate, loss function and relevant gradients.
+During step $t$, each worker $k$ first applies the update:
+$\mathbf{w}^{(k)}_{t}=\textsc{Optimize}(\mathbf{w}_{t-1}^{(k)}\,,\,\mathcal{B}_{t}^{(k)})$
+Moreover, operation $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$ computes and returns the average model vector (Li et al., 2020):
+$\overline{\mathbf{w}}_{t}=\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$
+In prior work, workers may (or not) synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}=\overline{\mathbf{w}}_{t}$.
+If synchronization is not performed at step $t$, then each worker continues with the locally updated model.
+Table 1. Notation
+Symbol Meaning
+[…]
+$t$ Time step index
+$K$ Number of workers
+$d$ Model dimension
+$\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$ Model of worker $k$
+$\mathbf{w}_{t}=[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}]$ Tensor of local models
+$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model vector
+$\overline{\mathbf{w}}_{t_{0}}$ Model after most recent sync.
+$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$ Local model drift
+$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ Model variance
+$\Theta$ Model variance threshold
+$\mathbf{\sigma}_{t}^{(k)}$ Lower-dim. proxy of $\mathbf{u}_{t}^{(k)}$
+$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)$ State of worker $k$
+$\overline{\mathbf{S}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{S}_{t}^{(k)}$ Average state
+Model Variance and FDA. The model variance quantifies the dispersion or spread of worker models around the average model:
+(2) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t}\right\|_{2}^{2}$
 This measure provides insight into how closely aligned the workers’ models are at any given time.
 High variance indicates that the models are widely spread out, essentially drifting apart, leading to a lack of cohesion in the aggregated model.
-Conversely, a moderate-low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective.
-The variance plays a critical role in our approach, as it helps us gauge the state of the training process making informed decisions about round termination.
-Round Invariant. We introduce the Round Invariant (RI), a condition that bounds model variance below a threshold $\Theta$:
+Conversely, a moderate or low variance suggests that the workers’ models are closely aligned, working collectively towards the shared objective.
+The FDA algorithm (Algorithm 1) is based on the premise that, as long as the variance is below a threshold $\Theta$, synchronization is not needed. Thus, we introduce the Round Invariant (RI):
 (3) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
-Our algorithm continuously monitors the model variance to maintain the RI.
-If the variance estimate exceeds the threshold $\Theta$ at any point $t$, the training round immediately terminates.
-This triggers synchronization, where all local model weights $\mathbf{w}^{(k)}_{t}$ are averaged and consolidated into the new global model $\overline{\mathbf{w}}_{t}$.
-Monitoring the RI.
-At this point, our focus shifts to devising a method to monitor the RI, as defined in (4).
-We begin by defining the local state for each worker $k$, denoted by $S_{k}(t)\in\mathbb{R}^{p}$, which encapsulates information from worker $k$ at time $t$; it is updated arbitrarily.
-Next, we introduce the global state, $S(t)\in\mathbb{R}^{p}$, which represents the collective state of the distributed system at time $t$.
-(5) $S(t)=\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)$
-Lastly, we define a special non-linear function $H:\mathbb{R}^{p}\to\mathbb{R}$, with the following property:
-(6) $\;H(S(t))\leq\Theta\;\Rightarrow\;\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta\quad\forall\;t$
-Conceptually, if we define $S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}\times\mathbb{R}^{d}$ and $H(v,x)=v-\left\|x\right\|_{2}^{2}$, then, it immediately follows from (3), (5) that $H\left(S\left(t\right)\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ and property (6) is satisfied.
-Direct monitoring of the RI is hindered by the high dimensionality of $S_{k}(t)$, with $d$ potentially ranging from millions to many billions, values commonly found in DNNs.
-To mitigate this communication burden, we must apply dimensionality reduction to the local drifts $\mathbf{u}_{t}^{(k)}$ and identify an appropriate function $H$, accordingly.
-However, this reduction in information within local states causes RI monitoring to become approximate.
-Upcoming sections will detail two strategies–termed ”linear”, and ”sketch”–each offering a different balance between communication efficiency and approximation accuracy.
-Algorithm. The proposed FDA algorithm is formalized in Algorithm 1. To streamline notation, $t$ also denotes each training round, which unfolds into three main parts:
-  1. (1)
-    Broadcast: Model $\mathbf{w}_{t}$ is distributed to all workers (Line 3).
-  2. (2)
-    Local Training: As long as we can guarantee the RI, workers continue training concurrently (Lines 4-9).
-Specifically, workers perform in-parallel updates to their local models $\mathbf{w}^{(k)}_{t}$ using mini-batch SGD, or some other optimization algorithm, by processing their local data (Lines 5-7).
-After each update, workers compute and broadcast their local states $S_{k}(t)$ (Line 8).
-Then, the global state $S(t)$ is calculated via (5), and using the function $H$, an estimate of the model variance is derived and compared with the threshold $\Theta$ (Line 9).
-Depending on this comparison, the local training stage either repeats or terminates.
-  3. (3)
-    Model Averaging: This stage is triggered when we can no longer guarantee the RI, i.e., $H(S(t))>\Theta$.
-At this point, all local models are communicated and then averaged to form $\mathbf{w}_{t+1}$, that is, the next round’s global model (Lines 10-12).
+To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each node $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$.
+These state vectors are vital for the subsequent estimation of the model variance, and underpin the two variants of the FDA algorithm (provided in Sections 3.1 and 3.2, respectively).
+Our estimation techniques begin by performing AllReduce on the states $\mathbf{S}_{t}^{(k)}$, consolidating them into the average state $\overline{\mathbf{S}}_{t}$ (Line 7).
+Importantly, this communication step requires significantly less bandwidth and resources than transmitting the full models $\mathbf{w}^{(k)}_{t}$.
+For each FDA variant, we also define a (different) function $H(\overline{\mathbf{S}}_{t})$ that overestimates the variance, i.e., it ensures that as long as $H(\overline{\mathbf{S}}_{t})\leq\Theta$ then the variance is bounded by $\Theta$.
+This guarantee is probabilistic for the Sketch-based variant of FDA, and deterministic for its Linear counterpart.
+Consequently, if $H(\overline{\mathbf{S}}_{t})>\Theta$ then synchronization is performed (Lines 8-9) — the RI invariant cannot be guaranteed.
+After synchronization, the model variance is zero.
+Efficiently Monitoring the RI. Estimating model variance efficiently is at the heart of FDA. To this end, we first introduce the local model drift
+$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$
+where $\overline{\mathbf{w}}_{t_{0}}$ is the model vector after the most recent synchronization. Subsequently, the model variance can be written as:
+(4) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+Proof. Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
+$\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}-\overline{\mathbf{w}}_{t_{0}}\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{u}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\frac{1}{K}\sum_{k=1}^{K}\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-2\left\langle\mathbf{u}_{t}^{(k)}\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\right)$
+$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left(\frac{1}{K}\sum_{k=1}^{K}\left\langle\mathbf{u}_{t}^{(k)}\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right)+\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\right)$
+$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\langle\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\langle\overline{\mathbf{u}}_{t}\,,\,\overline{\mathbf{u}}_{t}\right\rangle+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-2\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}+\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
 Algorithm 1 Federated Dynamic Averaging - FDA
-1:Require: $S_{k}(t)$: The local state $S_{k}(t)\in\mathbb{R}\times\mathbb{R}^{p}$ with $p\ll d$
-2:Require: $H(x)$: A function s.t. $H(S(t))\leq\Theta\Rightarrow\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
-3:Require: $K$: The number of workers indexed by $k$
-4:Require: $\Theta$: The model variance threshold
-5:Require: $b$: The local mini-batch size
-6:
-7:Initialize $\mathbf{w}_{1}\in\mathbb{R}^{d}$
-8:for each round $t=1,2,\dots$ do
-9:     communicate $\mathbf{w}_{t}$ to all workers
-10:     repeat
-11:          for each worker $k=1,\dots,K$ in parallel do
-12:               $\mathcal{B}_{k}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
-13:               $\mathbf{w}^{(k)}_{t}\leftarrow\mathbf{w}^{(k)}_{t}-\eta\nabla\ell(\mathbf{w}^{(k)}_{t};\mathcal{B}_{k})$triangleright In-place
-14:          communicate $S_{k}(t)$
-15:     until $H(S(t))>\Theta$triangleright As per (5)
-16:     for each worker $k=1,\dots,K$ in parallel do
-17:          communicate $\mathbf{w}^{(k)}_{t}$
-18:     $\mathbf{w}_{t+1}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}^{(k)}_{t}$
-19:
-3.1. Linear Approximation
-[…]
-$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
-$H\left(v\,,\,x\right)=v-x^{2}$
-Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$, since:
-$H\left(S\left(t\right)\right)$ =(5)${=}H\left(\frac{1}{K}\sum_{k=1}^{K}S_{k}(t)\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{k=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
-$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
-Furthermore, a random choice of $\xi$ is likely to perform poorly (terminate a round prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$.
-A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$.
-A heuristic choice is to take $\overline{\mathbf{u}}_{t_{0}}$, i.e., the change vector right before the current round started.
-All workers can estimate this without communication cost as the difference of the models at the beginning of the current and previous rounds and use $\xi=\frac{w_{t_{0}}-w_{t_{-1}}}{\left\|w_{t_{0}}-w_{t_{-1}}\right\|_{2}}.$
-3.2. Sketch Approximation
-A better estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005).
-An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix $\Xi$:
-$\operatorname{sk}\left(\mathbf{v}\right)=\Xi=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
+1:Require: $K$: The number of workers indexed by $k$
+2:Require: $\Theta$: The model variance threshold
+3:Require: $b$: The local mini-batch size
+4:
+5:Initialize $\mathbf{w}_{0}^{(k)}=\overline{\mathbf{w}}_{0}\in\mathbb{R}^{d}$
+6:for each step $t=1,2,\dots$ do
+7:     for each worker $k=1,\dots,K$ in parallel do
+8:          $\mathcal{B}_{t}^{(k)}\leftarrow$ (sample a batch of size $b$ from $\mathcal{D}_{k}$)
+9:          $\mathbf{w}^{(k)}_{t}\leftarrow\textsc{Optimize}(\mathbf{w}_{t-1}^{(k)}\,,\,\mathcal{B}_{t}^{(k)})$
+10:          Update $S_{t}^{(k)}$
+11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(S_{t}^{(k)})$
+12:          if $H(\overline{\mathbf{S}}_{t})>\Theta$ then
+13:               $\mathbf{w}_{t}^{(k)}\leftarrow\textsc{AllReduce}(\mathbf{w}^{(k)}_{t})$triangleright In-place
+Equation (4) indicates that the model variance estimation is reduced to estimating $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ efficiently.
+Computing $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ exactly can be done by an AllReduce operation, but this would incur a communication cost equal to performing synchronization, since the dimension of model drifts is equal to that of the models.
+We next present two techniques for communication-efficient variance estimation.
+To present them in a uniform manner, we define the local state $\mathbf{S}_{t}^{(k)}$, and the average state $\overline{\mathbf{S}}_{t}$ as
+$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)\quad,\quad\overline{\mathbf{S}}_{t}=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\frac{1}{K}\sum_{k=1}^{K}\mathbf{\sigma}_{t}^{(k)}\right)$
+where vectors $\mathbf{\sigma}_{t}^{(k)}$ are lower-dimensional representations of $\mathbf{u}_{t}^{(k)}$.
+In the two subsequent sections, we formally define $\mathbf{\sigma}_{t}^{(k)}$, and a function $H$ that is applied on the average state vector $\overline{\mathbf{S}}_{t}$.
+Each specification of $\mathbf{\sigma}_{t}^{(k)}$ and $H$ corresponds to a different FDA variant.
+3.1. SketchFDA: Sketch-based Estimation
+An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005).
+An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix:
+$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
 The $\operatorname{sk}(\cdot)$ operator is linear and can be computed in $\mathcal{O}\left(l\cdot d\right)$ steps.
@@
 The function
-$\mathcal{M}_{2}\left(\Xi\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
+$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
 is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$.
-We then define:
-$S_{k}(t)=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
-$H\left(\mathbf{v}\,,\,\Xi\right)=\mathbf{v}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\Xi\right)$
-Then, the condition $H\left(S\left(t\right)\right)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least $\left(1-\delta\right)$.
-To prove this, we proceed as follows: we first observe $\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)=\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)\leq(1+\epsilon)\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ with probability at least $(1-\delta)$.
-Then:
-$H\left(S\left(t\right)\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
-$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\;\text{with prob.
-at least}\;(1-\delta)$
+In this approach, the lower-dimensional representation of the local drifts becomes $\sigma_{t}^{(k)}=\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}^{l\times m}$.
+Theorem 3.1.
+Define:
+$S_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
+$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right).$
+Then, the condition $H\large(\overline{\mathbf{S}}_{t}\large)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$ with probability at least ($1-\delta$).
+Proof.
+$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{i=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)$
+${=}\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\operatorname{sk}\left(\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right)\right)$
+$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\operatorname{sk}\left(\overline{\mathbf{u}}_{t}\right)\right)$
+${\geq}\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}\quad\text{with prob. at least ($1-\delta$)}$
+$=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
+We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ with probability at least ($1-\delta$), i.e., we overestimate the model variance with probability at least ($1-\delta$), completing the proof.
+3.2. LinearFDA: Linear Approximation
+Although AMS sketches provide good estimates for variance, their dimension is in the several hundreds, and the communication cost of AllReduce on sketches, performed at each step, may be non-negligible.
+Therefore, we also examine a minimum-cost, ad-hoc estimation variant.
+In this approach, $\mathbf{\sigma}_{t}^{(k)}$ is reduced to a scalar, that is, $\sigma_{t}^{(k)}=\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\in\mathbb{R}$ where $\xi\in\mathbb{R}^{d}$ is any unit vector.
+Theorem 3.2.
+Define:
+$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
+$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
+Then, the condition $H\large(\overline{\mathbf{S}}_{t}\large)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$.
+Proof.
+$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
+$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
+$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}$
+$\geq\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\xi\right\|_{2}^{2}\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
+$=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
+We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof.
+A random choice of $\xi$ is likely to perform poorly (force synchronization prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$.
+A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$.
+A heuristic choice is to take the difference of the models of the last two synchronizations, assuming that the current model drift is correlated with the model drift at the last synchronization.
+All workers can compute this heuristic choice of $\xi$ without communication cost:
+$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}.$
+3.3. Discussion
+The main intuition for FDA is summarized in making the decision to synchronize dynamic, based on model variance during training.
+By Eq.
+4, model variance is the difference between the average of squared (local) drifts $\frac{1}{K}\sum\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}$ and the squared average (global) drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$.
+When both are low, or when both are high, synchronizing the local models is not necessary.
+When contrasting SketchFDA and LinearFDA, it should first be noted that both methods overestimate the true model variance.
+However, SketchFDA is probabilistically quite accurate.
+Our findings suggest that setting $l=5$ and $m=250$ (as defined in Section 3.1) yields excellent accuracy in practice.
+These values have been adopted in our experiments and are thus recommended.
+With these settings, the size of a sketch is $4k$ bytes, significantly smaller than the size of all our models.
+Between the two variants of FDA, LinearFDA requires less computational effort and bandwidth to compute and to transmit its local state (the norm of a vector plus the dot product of two vectors, rather than a sketch).
+On the other hand, SketchFDA provides in practice a very accurate estimator of the true model variance which is expected to lead to fewer synchronizations.
+An alternative point of view of FDA, relates to the well-known “bias vs.
+variance” trade-off (Goodfellow et al., 2016), which encapsulates the tension between under-fitting and over-fitting.
+Note that the term “variance” here refers to change to the learned model $\mathbf{w}$ when the training data changes, and is different from our model variance $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$.
 4. Experiments
@@
 Platform.
-We employ TensorFlow (Abadi et al., 2015), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments.
+We employ TensorFlow (Abadi et al., 2016), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments.
 We modified TensorFlow to include our FDA variants and all competitive algorithms.
+All relevant code, figures, and data of this study are available in https://github.com/miketheologitis/FedL-Sync-FDA.
 Table 2. Summary of Experiments
@@
 The core experiments involve training Convolutional Neural Networks (CNNs) of varying sizes and complexities on two datasets: MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2012).
-For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2014), denoted as VGG16*, consisting of 2.6 million parameters.
+For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2015), denoted as VGG16*, consisting of 2.6 million parameters.
 VGG16* was specifically adapted for the MNIST dataset, a less demanding learning problem compared to ImageNet (Russakovsky et al., 2015), for which VGG16 was designed.
@@
 Both models use Glorot uniform initialization (Glorot and Bengio, 2010).
-For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2016), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2016).
+For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2017), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2017).
 The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
-Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021; Ridnik et al., 2021).
-For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2020).
+Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021).
+For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2021).
 We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015).
-Following the feature extraction step (Goodfellow et al., 2016; Zhuang et al., 2019), the testing accuracy on CIFAR-100 stands at $60\%$.
-Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Yosinski et al., 2014).
+Following the feature extraction step (Goodfellow et al., 2016), the testing accuracy on CIFAR-100 stands at $60\%$.
+Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Nakandala and Kumar, 2022).
 Algorithms.
 We consider five distributed deep learning algorithms: LinearFDA, SketchFDA, Synchronous  [^2] The name was derived from the Bulk Synchronous Parallel approach; can be understood as a special case of the FDA Algorithm 1 where $\Theta$ is set to zero., FedAdam (Reddi et al., 2021), and FedAvgM (Hsu et al., 2019); the first three are standard in all experiments.
-Depending on the local optimizer, Adam (Kingma and Ba, 2017) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their federated counterparts FedAdam or FedAvgM, respectively.
+Depending on the local optimizer, Adam (Kingma and Ba, 2015) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their communication-efficient federated counterparts FedAdam or FedAvgM, respectively.
 Evaluation Methodology.
@@
 For experiments involving FedAvgM and FedAdam, we use $E=1$ local epochs, following (Reddi et al., 2021).
-For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2017).
+For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2015).
 In these cases, FedAdam also adheres to the default settings for both local and server optimization (Reddi et al., 2021; Chollet et al., 2015).
-For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2016).
+For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2017).
 For FedAvgM, local optimization is conducted with default settings (Hsu et al., 2019; Chollet et al., 2015), while server optimization employs SGD with momentum, setting the momentum parameter and learning rate to $0.9$ and $0.316$, respectively (Reddi et al., 2021).
-Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2017), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
+Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2019), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
 Data Distribution. In all experiments, the training dataset is divided into approximately equal parts among the workers. To assess the impact of data heterogeneity, we explore three scenarios:
@@
     Non-IID: Label $Y$ — All samples from label $Y$ are assigned to a few workers, while the rest are distributed in an IID manner.
-Main Findings.
-The main findings of our experimental analyses are: (1) LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance.
-(2) LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation.
-(3) The performance of LinearFDA and SketchFDA is comparable in most experiments.
-SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to a number).
+Figure 2.
+LeNet-5 on MNIST.
+At Non-IID: Label ”0”, the samples of Label ”0” are assigned to few workers.
+At Non-IID: 60%, 60% of the dataset is sorted and allocated to workers, causing some workers to receive many samples from the same label
+Figure 3. VGG16* on MNIST
+4.2. Main Findings
+The main findings of our experimental analyses are:
+  1. (1)
+    LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance.
+  2. (2)
+    LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation.
+  3. (3)
+    The performance of LinearFDA and SketchFDA is comparable in most experiments.
+SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to two numbers).
 SketchFDA significantly outperforms LinearFDA at the transfer learning scenario.
-(4) The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
+  4. (4)
+    The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
 4.3. Results
@@
 In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 2) that attained the Accuracy Target of $0.985$.
-Figure 1. LeNet-5 on MNIST
-Figure 2. VGG16* on MNIST
 Figure 4. DenseNet121 on CIFAR-10
 Figure 5. DenseNet201 on CIFAR-10
+Figure 6.
+Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right).
+Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively.
+A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model.
+Figure 7. LeNet-5 on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.98$
+Figure 8. VGG16* on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.994$
+Figure 9. DenseNet121 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.8$
+Figure 10. DenseNet201 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.78$
 FDA balances Communication vs.
@@
 Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost).
-This is evident in Figures 12,  3, and 4 (where Synchronous appears in the bottom right — low computation, very high communication).
+This is evident in Figures 234, and 5 (where Synchronous appears in the bottom right — low computation, very high communication).
 Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation.
-Indeed, as shown in Figures 1-4, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation.
-Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 1-4.
+Indeed, as shown in Figures 2-5, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation.
+Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 23,  4, and 5.
 In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency.
@@
 On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement.
-For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  34), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201).
+For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  4, and 5), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201).
 In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
@@
 The patterns of performance we highlighted are consistent across most of the conducted tests.
-Figure 5.
-Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right).
-Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively.
-A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
 Dependence on $K$.
@@
 In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup.
-This unpredictability is starkly illustrated with VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies.
-Specifically, Figure 6 demonstrates (at its upper part) that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication (shown at the upper part of the figure).
+This unpredictability is starkly illustrated with LeNet-5 and VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies.
+Figures 7, and 8 (left) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication.
 These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources.
-However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (upper part of Figures 78), the expected behavior starts to emerge.
-Specifically, for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies.
+However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (left part of Figures 910), the expected behavior starts to emerge.
+Especially for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies.
 Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead.
 Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers.
+Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 5-7.
+Specifically, they require up to 30 times less communication than FedAdam, 4 times less than FedAvgM, and up to 2.5 orders of magnitude less than Synchronous.
 FDA: Dependence on $\Theta$.
@@
 A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence.
-This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 67 and 8).
-Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 7 and 8.
+This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 7-10).
+Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 9 and 10.
 It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
-Figure 6. VGG16* on MNIST — Accuracy Target: $0.994$
-Figure 7. DenseNet121 on CIFAR-10 — Accuracy Target: $0.8$
-Figure 8. DenseNet201 on CIFAR-10 — Accuracy Target: $0.78$
+Figure 11. Empirical Estimation of the Variance Threshold
+Figure 12. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
 FDA: Choice of $\Theta$.
-The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 67, and 8.
+The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 7,  89, and 10.
 Therefore, identifying this range becomes crucial.
@@
 Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 2), thereby investigating different computation and communication trade-offs.
-For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by computation cost, rendering communication concerns negligible.
+For instance, in the ARIS-HPC environment (https://www.hpc.grnet.gr/en/hardware-2/) with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible.
 In such contexts, lower $\Theta$ values are favored due to their computational efficiency.
 On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
-To assist researchers in selecting an optimal variance threshold, Figure 9 presents empirical estimations for $\Theta$ across three distinct learning settings: FL (assuming a common channel of $0.5$Gbps), Balanced (communication-computation equilibrium), and HPC.
+To assist researchers in selecting the variance threshold, Figure 11 presents empirical estimations for $\Theta$ across three distinct learning settings: (1) FL (assuming a common channel of $0.5$Gbps), (2) Balanced communication-computation equilibrium setting, and (3) our HPC environment at the ARIS supercomputer.
 $\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
@@
 $\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
-Figure 9. Empirical Estimation of the Variance Threshold
 FDA: Linear vs.
 Sketch.
-In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 123, and  4.
-This suggests that the precision of the variance approximation is not critical; occasional ”unnecessary” synchronizations do not significantly impact overall performance.
+In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 234, and 5.
+This suggests that the precision of the variance approximation is not critical.
 However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 12).
 In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation.
-On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, particularly if communication-efficiency is paramount.
-Figure 10. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
+On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, if communication-efficiency is paramount.
 5. Conclusions
@@
 Our results push the limits of modern communication-efficient distributed deep learning, paving the way for scalable, dynamic, and broadly applicable strategies.
+Acknowledgements.
+This work was supported by the EU project CREXDATA under Horizon Europe agreement No.
+101092749.
+Moreover, this work was supported by computational time granted from the National Infrastructures for Research and Technology S.A.
+(GRNET S.A.) in the National HPC facility - ARIS - under project ID pa230902-fda1.

### v4

@@
 Abstract.
-Driven by the ever-growing volume and decentralized nature of data, coupled with the need to harness this data and generate knowledge from it, has led to the extensive use of distributed deep learning (DDL) techniques for training.
-These techniques rely on local training that is performed at the distributed nodes based on locally collected data, followed by a periodic synchronization process that combines these models to create a global model.
-However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability.
-Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules.
-These drawbacks also have a direct impact on the time required for the training process, necessitating excessive time for data communication.
+The ever-growing volume and decentralized nature of data, coupled with the need to harness it and extract knowledge, have led to the extensive use of distributed deep learning (DDL) techniques for training.
+These techniques rely on local training performed at distributed nodes using locally collected data, followed by a periodic synchronization process that combines these models to create a unified global model.
+However, the frequent synchronization of deep learning models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability.
+Worse yet, DDL algorithms typically waste valuable bandwidth and render themselves less practical in bandwidth-constrained federated settings by relying on overly simplistic, periodic, and rigid synchronization schedules.
+These inefficiencies make the training process increasingly impractical as they demand excessive time for data communication.
 To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance.
-In essence, the costly synchronization step is triggered only if the local models, which are initialized from a common global model after each synchronization, have significantly diverged.
-This decision is facilitated by the communication of a small local state from each distributed node/worker.
+In essence, the costly synchronization step is triggered only if the local models—initialized from a common global model after each synchronization—have significantly diverged.
+This decision is facilitated by the transmission of a small local state from each distributed node.
 Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms.
 Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
-Federated Learning, Distributed Learning
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 26th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Parallel computing methodologies
+copyright: rightsretaineddoi: isbn: 978-3-89318-098-1conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
 1. Introduction
 The big data era has been marked by an unprecedented scale of training datasets (Zhang et al., 2021; Rauf et al., 2024).
-These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations (especially in federated settings, such as drones collecting and collaboratively building a global model/view of an area), and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024).
-Such constraints complicate the use of Deep Learning (DL) techniques in the aforementioned scenarios, and provide challenges in harnessing knowledge from distributed data.
-Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024).
-Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local models with the global one.
-The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010).
-Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Miao et al., 2021; Fu et al., 2022) but compromise convergence speed and model quality.
-A significant challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Tang et al., 2023; Wang et al., 2016).
-This leads to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022).
-Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
+These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations—particularly in federated settings, such as drones collecting and collaboratively building a global model or view of an area—and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024).
+Such constraints pose significant challenges to the application of Deep Learning (DL) techniques.
+Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient learning over large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024).
+Most DDL methods are iterative, where each iteration involves a phase of local training followed by the synchronization of the local models with the global one.
+The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), averages the local model updates and applies the averaged update to each local model (Zinkevich et al., 2010).
+Less synchronized variants have also been proposed to ameliorate the effect of straggler workers, but they often compromise convergence speed and model quality (Miao et al., 2021; Fu et al., 2022).
+A significant challenge inherent in the traditional techniques, especially in federated DL settings where models are huge and worker interconnections are slow, is the communication bottleneck, which restricts system scalability (Tang et al., 2023; Wang et al., 2016).
+Specifically, the communication bottleneck arises from the frequent exchange (synchronization) of model parameters—often in the range of billions—across distributed workers.
+The synchronization process entails substantial data volume transfer and generally dominates the overall training time, leading to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022).
+Addressing this challenge to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
 The most direct method to alleviate the communication burden is to reduce the frequency of communication rounds.
@@
 To further reduce communication costs of Local-SGD, more sophisticated communication strategies introduce varying sequences of local update steps $\{\tau_{0},...,\tau_{R}\}$, instead of a fixed $\tau$.
-In (Wang and Joshi, [n.
-d.]), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps.
+In (Wang and Joshi, [n.d.]), in order to minimize convergence error with respect to wall-time, the authors proposed a decreasing sequence of local update steps.
 Conversely, the focus in (Haddadpour et al., 2019) was on reducing the number of communication rounds for a fixed number of model updates and an increasing sequence emerged.
@@
 FDA utilizes a novel 2-action, conditional synchronization protocol, designed to avoid the need to decide or guess the proper values of local update steps, or to synchronize after each training step, but rather only performs the costly synchronization process when needed.
-Our FDA algorithm dynamically triggers synchronization based on the value of model variance across nodes.
+Our FDA algorithm dynamically triggers synchronization based on the value of model variance across worker-nodes.
 In a nutshell, the costly synchronization step is only triggered if the local models have diverged significantly, which implies that the global model may no longer be accurate.
-As Figure 1 demonstrates, at the start, nodes enter the local training step with the same global model (Figure 1.(A)).
-Then, local training commences and each distributed node/worker computes its local state (Figure 1.(B)).
-This is followed by the transmission (Figure 1.(C)) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size.
-During transmission, the local states are aggregated and their average is made available to all nodes—an operation known as AllReduce.
+As Figure 1 demonstrates, at the start, workers enter the local training step with the same global model (Figure 1.A).
+Then, local training commences and each distributed worker-node computes its local state, which encapsulates helpful information for estimating the model variance (Figure 1.B).
+This is followed by the transmission (Figure 1.C) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size.
+During transmission, the local states are aggregated and their average is made available to all workers—an operation known as AllReduce.
 This operation does not require (or prohibit) the use of a central node.
-Based on the aggregated state, the nodes can estimate (Figure 1.(D)) whether the variance of the local models may have exceeded a threshold.
-If this is not the case, the costly synchronization step (Figure 1.(E)) is avoided and local training continues.
-What is important is how to properly pick these local states computed at, and then transmitted by, the local nodes.
+Based on the aggregated state, the workers can estimate (Figure 1.D) whether the variance of the local models may have exceeded a threshold.
+If this is not the case, the costly synchronization step (Figure 1.E) is avoided and local training continues.
+What is important is how to properly pick these local states computed at, and then transmitted by, the local workers.
 To address this problem, we propose two variants of our FDA algorithm.
 Our contributions can be summarized as follows:
-  • We propose FDA, an algorithm that dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold.
-This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
-  • We propose two variants of FDA, which differ on the local state that is computed at, and then transmitted by, each distributed node/worker. These two variants are termed SketchFDA and LinearFDA.
+  • We propose FDA, an algorithm that dynamically decides to synchronize local workers when model variance across workers exceeds a threshold.
+This strategy drastically reduces communication, while preserving cohesive progress towards the shared training objective.
+  • We propose two variants of FDA, which differ in the amount of information preserved in the local states that are transmitted by each worker and aggregated for subsequent estimation of model variance.
+These two variants, termed SketchFDA and LinearFDA, offer a different balance between communication efficiency and approximation accuracy.
   • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks.
-Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance.
+Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by 1-2 orders of magnitude in communication savings, while maintaining equivalent model performance.
 Furthermore, it effectively balances the competing demands of communication and computation, providing greatly improved trade-offs.
-  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case.
-We note that in Federated Learning scenarios, the norm is that data are not independently and identically distributed across nodes.
+  • We demonstrate FDA’s robustness in various challenging Non-IID settings, common in real-world Federated Learning applications.
+While state-of-the-art methods typically require substantially more resources to converge under Non-IID conditions, FDA maintains consistent and comparable performance across both IID and Non-IID settings.
 Outline.
@@
 FDA.
-The local training step is followed by the computation of a local state by all nodes.
+The local training step is followed by the computation of a local state by all worker-nodes.
 Then, the (small in size) local states are aggregated.
-Based on the aggregated result, all nodes estimate if synchronization is required.
+Based on the aggregated result, all workers estimate if synchronization is required.
 In most cases, the expensive synchronization step of the models is avoided and local training continues
@@
 Consider distributed training of deep neural networks over multiple workers (Dean et al., 2012; Li et al., 2021).
-In this setting, each worker has access to its own set of training data $\mathcal{D}_{k}$, and the collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss.
+In this setting, each worker represents a data owner (equivalently, a local model owner) and has access to its own set of training data $\mathcal{D}_{k}$.
+Workers can utilize any available hardware they possess (e.g., GPUs, CPUs) to perform learning steps.
+The collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss.
 This scenario can be effectively modeled as a distributed optimization problem, formulated as follows:
@@
 Function $\ell(\mathbf{w};\zeta_{k})$ represents the loss for data sample $\zeta_{k}$ given model $\mathbf{w}$.
+Solution direction.
+As noted in the seminal work (Kairouz et al., 2021), research in FL should focus primarily on synchronous solutions.
+This allows different lines of research (e.g., compression, privacy, etc.) to be developed independently and then combined seamlessly.
+Our work, along with most communication-efficient FL strategies, adheres to this synchronous paradigm.
+However, such approaches may be less effective in environments where each communication operation incurs significant overhead regardless of the size of the data being transmitted (e.g., high-latency).
+In these scenarios, asynchronous mechanisms become necessary, though they typically fall outside the primary focus of contemporary FL research.
+That said, FDA can be modified to work asynchronously (as explained in Section 3.3).
 Communication efficient Local-SGD.
 The work in (Li et al., 2021) decomposes each round into two phases.
-In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds.
-In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives.
+In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase runs $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds.
+In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives and fixed local update steps for other types of local objectives.
 The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD.
@@
 In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
-Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021).
+Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative of communication efficient Local-SGD algorithms, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021).
 In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints.
 Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
+Unlike previous approaches that rely on predetermined synchronization schedules (fixed, decaying, or otherwise), our work introduces a dynamic synchronization strategy.
+FDA adapts continuously during the training process, basing synchronization decisions on a real-time metric: the model variance across workers.
 Accelerating convergence.
-An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence.
-Consequently, recent works have deployed accelerated versions of SGD to the distributed setting.
+An indirect, yet highly effective way to mitigate the communication burden in DDL, is to speed up convergence.
+Consequently, recent works have built upon communication efficient Local-SGD methods by deploying accelerated versions of SGD to the distributed setting.
 Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2015) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a).
@@
 Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
+While these approaches primarily focus on enhancing the optimization process and typically employ fixed synchronization intervals (e.g., every local epoch), our work addresses a complementary aspect: determining the optimal timing for synchronization.
+FDA’s dynamic synchronization strategy is orthogonal to these optimization techniques and can be integrated with them by simply adjusting the synchronization decision.
 Compression.
@@
 Key strategies include sparsification, where only crucial components of information are transmitted, as explored in (Aji and Heafield, 2017), and quantization techniques, which involve transmitting only quantized gradients, as detailed in (Shlezinger et al., 2021).
-Notably, these compression techniques can be combined with Local-SGD methods to enhance communication-efficiency further.
-A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies.
+These techniques can be combined with Local-SGD methods to enhance communication-efficiency further.
+An example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings.
+Crucially, FDA is fully compatible with any technique that reduces the cost of synchronization (e.g.
+model compression).
+Our approach simply adjusts the timing of the synchronization decision without altering the data being synchronized.
+This ensures that any compression technique effective in traditional methods (BSP, Local-SGD, etc.) will be equally effective when deployed with FDA.
+Therefore, the communication savings demonstrated in the relevant literature (Wang et al., 2023) can be safely expected to carry over to our approach as well.
 Additionally, sketching emerges as another fundamental tool in large-scale machine learning.
@@
 $\mathbf{w}_{t}=\left[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}\right]\quad,\quad\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$
-Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data.
+Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data.
 It incorporates the learning rate, loss function and relevant gradients.
@@
 $\overline{\mathbf{w}}_{t}=\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$
-In prior work, workers may (or not) synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}=\overline{\mathbf{w}}_{t}$.
-If synchronization is not performed at step $t$, then each worker continues with the locally updated model.
+Workers synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}:=\overline{\mathbf{w}}_{t}$.
+If synchronization is not performed at step $t$, each worker continues training with its locally updated model.
+A comprehensive list of the notation used throughout this section is provided in Table 1.
 Table 1. Notation
@@
 $d$ Model dimension
+$\mathcal{D}_{k}$ Training data of worker $k$
+$\mathcal{B}^{(k)}_{t}$ A batch sampled from $\mathcal{D}_{k}$
 $\mathbf{w}_{t}^{(k)}\in\mathbb{R}^{d}$ Model of worker $k$
 $\mathbf{w}_{t}=[\mathbf{w}_{t}^{(1)},\ldots,\mathbf{w}_{t}^{(K)}]$ Tensor of local models
-$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model vector
+$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model (global model)
 $\overline{\mathbf{w}}_{t_{0}}$ Model after most recent sync.
+$\overline{\mathbf{w}}_{t_{-1}}$ Model after 2nd most recent sync.
 $\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$ Local model drift
+$\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}$ Average model drift (global drift)
 $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ Model variance
 $\Theta$ Model variance threshold
-$\mathbf{\sigma}_{t}^{(k)}$ Lower-dim. proxy of $\mathbf{u}_{t}^{(k)}$
-$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)$ State of worker $k$
+$\mathbf{S}_{t}^{(k)}$ State of worker $k$
 $\overline{\mathbf{S}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{S}_{t}^{(k)}$ Average state
+$H(\cdot)$ Function for variance estimation
+$\operatorname{sk}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{l\times m}$ AMS sketch operator (§3.1)
+$\mathcal{M}_{2}(\cdot):\mathbb{R}^{l\times m}\to\mathbb{R}$ $L^{2}$ norm squared estimate (§3.1)
+$\epsilon$ Error of sketch estimate (§3.1)
+$(1-\delta)$ Confidence of approximation (§3.1)
+$l=\mathcal{O}(\log 1/\delta)$Rows of sketch matrix (§3.1)
+$m=\mathcal{O}(1/\epsilon^{2})$Columns of sketch matrix (§3.1)
+$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}$ Heuristic vec. for LinearFDA3.2)
 Model Variance and FDA. The model variance quantifies the dispersion or spread of worker models around the average model:
@@
 (3) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$
-To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each node $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$.
+To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each worker $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$.
 These state vectors are vital for the subsequent estimation of the model variance, and underpin the two variants of the FDA algorithm (provided in Sections 3.1 and 3.2, respectively).
@@
 After synchronization, the model variance is zero.
-Efficiently Monitoring the RI. Estimating model variance efficiently is at the heart of FDA. To this end, we first introduce the local model drift
-$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$
-where $\overline{\mathbf{w}}_{t_{0}}$ is the model vector after the most recent synchronization. Subsequently, the model variance can be written as:
+Efficiently Monitoring the RI.
+Estimating model variance efficiently is at the heart of FDA.
+To this end, we first introduce the local model drift, $\mathbf{u}_{t}^{(k)}$, and average drift, $\overline{\mathbf{u}}_{t}$, defined as follows:
+$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}\quad,\quad\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}$
+Here, $\overline{\mathbf{w}}_{t_{0}}$ denotes the model vector after the most recent synchronization. Subsequently, the model variance can be written as:
 (4) $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}^{(k)}_{t}\right\|_{2}^{2}\right)-\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
-Proof. Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
+Proof.
+Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
 $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ $=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}-\overline{\mathbf{w}}_{t_{0}}\right)=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{u}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}-\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$
@@
 9:          $\mathbf{w}^{(k)}_{t}\leftarrow\textsc{Optimize}(\mathbf{w}_{t-1}^{(k)}\,,\,\mathcal{B}_{t}^{(k)})$
-10:          Update $S_{t}^{(k)}$
-11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(S_{t}^{(k)})$
+10:          Update $\mathbf{S}_{t}^{(k)}$
+11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(\mathbf{S}_{t}^{(k)})$
 12:          if $H(\overline{\mathbf{S}}_{t})>\Theta$ then
 13:               $\mathbf{w}_{t}^{(k)}\leftarrow\textsc{AllReduce}(\mathbf{w}^{(k)}_{t})$triangleright In-place
-Equation (4) indicates that the model variance estimation is reduced to estimating $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ efficiently.
-Computing $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ exactly can be done by an AllReduce operation, but this would incur a communication cost equal to performing synchronization, since the dimension of model drifts is equal to that of the models.
-We next present two techniques for communication-efficient variance estimation.
-To present them in a uniform manner, we define the local state $\mathbf{S}_{t}^{(k)}$, and the average state $\overline{\mathbf{S}}_{t}$ as
-$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)\quad,\quad\overline{\mathbf{S}}_{t}=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\frac{1}{K}\sum_{k=1}^{K}\mathbf{\sigma}_{t}^{(k)}\right)$
-where vectors $\mathbf{\sigma}_{t}^{(k)}$ are lower-dimensional representations of $\mathbf{u}_{t}^{(k)}$.
-In the two subsequent sections, we formally define $\mathbf{\sigma}_{t}^{(k)}$, and a function $H$ that is applied on the average state vector $\overline{\mathbf{S}}_{t}$.
-Each specification of $\mathbf{\sigma}_{t}^{(k)}$ and $H$ corresponds to a different FDA variant.
+Conceptually, following Eq (4), to precisely monitor the variance, we need to calculate two quantities: (1) $\frac{1}{K}\sum_{k=1}^{K}\|\mathbf{u}^{(k)}_{t}\|_{2}^{2}$, and (2) $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$.
+The first quantity requires an AllReduce operation on the squared norm of the worker drifts, which involves minimal overhead since these values are scalar.
+In contrast, the second quantity necessitates an AllReduce operation on the worker drifts themselves, which are of model dimension, thus incurring a high communication cost.
+In fact, this operation is equivalent to synchronization, which is exactly what we aim to avoid in the first place.
+Thus, it becomes evident that communication-efficient model variance estimation hinges on estimating $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$ efficiently.
+Upcoming sections will detail two techniques for communication efficient variance estimation (which primarily involves estimating $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$): SketchFDA and LinearFDA.
+To present them uniformly, we introduce the local state $\mathbf{S}_{t}^{(k)}$, a tensor which contains: (1) the scalar value $\|\mathbf{u}^{(k)}_{t}\|_{2}^{2}$ for precisely calculating the first quantity, and (2) a low-dimensional summary of $\mathbf{u}_{t}^{(k)}$, different for each technique, for estimating the second quantity.
+For each technique we define an estimation function $H(\cdot)$ that calculates the current variance estimate from average state $\overline{\mathbf{S}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{S}_{t}^{(k)}$ (obtained via AllReduce).
 3.1. SketchFDA: Sketch-based Estimation
-An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization of AMS sketches, as detailed in (Cormode and Garofalakis, 2005).
+An optimal estimator for $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ can be obtained through the utilization and properties of AMS sketches, as detailed in (Cormode and Garofalakis, 2005).
 An AMS sketch of a vector $\mathbf{v}\in\mathbb{R}^{d}$ is an $l\times m$ real matrix:
-$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\xi_{1}&\xi_{2}&\dots&\xi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
-The $\operatorname{sk}(\cdot)$ operator is linear and can be computed in $\mathcal{O}\left(l\cdot d\right)$ steps.
-Let $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$.
-The function
-$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\xi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
-is an excellent estimator of $\left\|\mathbf{v}\right\|_{2}^{2}$, as with probability at least $\left(1-\delta\right)$, $\mathcal{M}_{2}\left(\operatorname{sk}\left(\mathbf{v}\right)\right)$ is within $\epsilon$-relative error of $\left\|\mathbf{v}\right\|_{2}^{2}$.
-In this approach, the lower-dimensional representation of the local drifts becomes $\sigma_{t}^{(k)}=\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\in\mathbb{R}^{l\times m}$.
+$\operatorname{sk}\left(\mathbf{v}\right)=\begin{bmatrix}\psi_{1}&\psi_{2}&\dots&\psi_{l}\end{bmatrix}^{\top}\in\mathbb{R}^{l\times m}\;\;,\;\;l\cdot m\ll d$
+An estimate for squared-norm $\left\|\mathbf{v}\right\|_{2}^{2}$ is provided by the formula
+$\mathcal{M}_{2}\left(\operatorname{sk}(\mathbf{v})\right)=\mathrm{median}\left\{\left\|\psi_{i}\right\|_{2}^{2}\;,\;i=1,\dots,l\right\}$
+The quality of estimation depends on the size of the sketch.
+For chosen $\epsilon,\delta>0$, where sketch dimensions are given by $l=\mathcal{O}\left(\log 1/\delta\right)$ and $m=\mathcal{O}\left(1/\epsilon^{2}\right)$, we have the following probabilistic guarantee: with confidence at least $1-\delta$,
+$\mathcal{M}_{2}(\operatorname{sk}(\mathbf{v}))\in(1\pm\epsilon)\left\|\mathbf{v}\right\|_{2}^{2}$
+Notably, observe that the accuracy ($\epsilon$) and confidence ($1-\delta$) only depend on the size of the sketch and not on the dimensionality of vector $\mathbf{v}$.
+Two crucial properties of the AMS sketch are that (a) it is a linear transformation, i.e., for $\alpha_{1},\alpha_{2}\in\mathbb{R}$ and $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{d}$,
+$\operatorname{sk}(\alpha_{1}\mathbf{v}_{1}+\alpha_{2}\mathbf{v}_{2})=\alpha_{1}\operatorname{sk}(\mathbf{v}_{1})+\alpha_{2}\operatorname{sk}(\mathbf{v}_{2})$
+and (b) can be computed efficiently in time $O(l\cdot d)$.
+In the SketchFDA approach, the salient idea is to employ AMS sketches $\operatorname{sk}(\mathbf{u}_{t}^{(k)})\in\mathbb{R}^{l\times m}$ as a low-dimensional representation of the local drifts $\mathbf{u}_{t}^{(k)}$.
 Theorem 3.1.
-Define:
-$S_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
+Let $l=\mathcal{O}(\log\frac{1}{\delta})$ and $m=\mathcal{O}(\frac{1}{\epsilon^{2}})$. Define the local state as
+$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right)\in\mathbb{R}\times\mathbb{R}^{l\times m}$
+and the approximation function as
 $H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\frac{1}{1+\epsilon}\mathcal{M}_{2}\left(\frac{1}{K}\sum_{k=1}^{K}\operatorname{sk}\left(\mathbf{u}_{t}^{(k)}\right)\right).$
@@
 We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$ with probability at least ($1-\delta$), i.e., we overestimate the model variance with probability at least ($1-\delta$), completing the proof.
+∎
+In Section 3.3, we discuss the empirical basis for choosing the values of $l$ and $m$, and how they practically impact the quality of the sketch approximation.
 3.2. LinearFDA: Linear Approximation
 Although AMS sketches provide good estimates for variance, their dimension is in the several hundreds, and the communication cost of AllReduce on sketches, performed at each step, may be non-negligible.
-Therefore, we also examine a minimum-cost, ad-hoc estimation variant.
-In this approach, $\mathbf{\sigma}_{t}^{(k)}$ is reduced to a scalar, that is, $\sigma_{t}^{(k)}=\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\in\mathbb{R}$ where $\xi\in\mathbb{R}^{d}$ is any unit vector.
+Therefore, we also introduce a low-cost, ad-hoc estimation variant.
+In this approach, instead of an AMS sketch, each local state contains the scalar value $\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\rangle\in\mathbb{R}$, where $\xi\in\mathbb{R}^{d}$ is a unit vector, known to all workers.
 Theorem 3.2.
-Define:
+Define the local state as
 $\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right)\in\mathbb{R}\times\mathbb{R}\;\;,\;\;\left\|\xi\right\|_{2}=1$
-$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
+and the approximation function as
+$H\left(\overline{\mathbf{S}}_{t}\right)=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
 Then, the condition $H\large(\overline{\mathbf{S}}_{t}\large)\leq\Theta$ implies $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)\leq\Theta$.
 Proof.
-$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
-$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left\|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right\|_{2}^{2}$
+$H\left(\overline{\mathbf{S}}_{t}\right)$ $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\frac{1}{K}\sum_{i=1}^{K}\left\langle\xi\,,\,\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
+$=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\frac{1}{K}\sum_{i=1}^{K}\mathbf{u}_{t}^{(k)}\right\rangle\right|^{2}$
 $=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}-\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}$
@@
 $=\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$
-We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof.
-A random choice of $\xi$ is likely to perform poorly (force synchronization prematurely), as it is likely to be close to orthogonal to $\overline{\mathbf{u}}_{t}$.
-A good choice would be a vector correlated to $\overline{\mathbf{u}}_{t}$.
-A heuristic choice is to take the difference of the models of the last two synchronizations, assuming that the current model drift is correlated with the model drift at the last synchronization.
-All workers can compute this heuristic choice of $\xi$ without communication cost:
-$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}.$
+We proved that $H(\overline{\mathbf{S}}_{t})\geq\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$, i.e., we always overestimate the model variance, completing the proof. ∎
+An arbitrary choice of $\xi$ (e.g., a random vector) is likely to estimate $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ poorly; if $\mathbf{\xi}$ is uncorrelated to $\overline{\mathbf{u}}_{t}$, then $\left|\left\langle\xi\,,\,\overline{\mathbf{u}}_{t}\right\rangle\right|^{2}$ will likely be close to zero.
+A heuristic choice that might be correlated to $\overline{\mathbf{u}}_{t}$ is the (normalized) value of $\overline{\mathbf{u}}_{t_{0}}$, the global drift vector right at the time of last synchronization.
+All nodes can compute it independently without extra communication, if they take the difference of the models of the last two synchronizations:
+$\mathbf{\xi}=\frac{\overline{\mathbf{u}}_{t_{0}}}{\left\|\overline{\mathbf{u}}_{t_{0}}\right\|_{2}}=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}$
 3.3. Discussion
+FDA: Intuition.
 The main intuition for FDA is summarized in making the decision to synchronize dynamic, based on model variance during training.
-By Eq.
-4, model variance is the difference between the average of squared (local) drifts $\frac{1}{K}\sum\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}$ and the squared average (global) drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$.
-When both are low, or when both are high, synchronizing the local models is not necessary.
-When contrasting SketchFDA and LinearFDA, it should first be noted that both methods overestimate the true model variance.
-However, SketchFDA is probabilistically quite accurate.
-Our findings suggest that setting $l=5$ and $m=250$ (as defined in Section 3.1) yields excellent accuracy in practice.
-These values have been adopted in our experiments and are thus recommended.
-With these settings, the size of a sketch is $4k$ bytes, significantly smaller than the size of all our models.
-Between the two variants of FDA, LinearFDA requires less computational effort and bandwidth to compute and to transmit its local state (the norm of a vector plus the dot product of two vectors, rather than a sketch).
-On the other hand, SketchFDA provides in practice a very accurate estimator of the true model variance which is expected to lead to fewer synchronizations.
-An alternative point of view of FDA, relates to the well-known “bias vs.
-variance” trade-off (Goodfellow et al., 2016), which encapsulates the tension between under-fitting and over-fitting.
-Note that the term “variance” here refers to change to the learned model $\mathbf{w}$ when the training data changes, and is different from our model variance $\operatorname{\mathcal{V}\!\text{{ar}}}\left(\mathbf{w}_{t}\right)$.
+This metric is designed to capture the collective state of the training process.
+In what follows, we provide intuition on why this is the case.
+It is important to remember that the global model $\overline{\mathbf{w}}_{t}$ and, by extension, the global drift $\overline{\mathbf{u}}_{t}$, are ultimately what we care about and evaluate.
+Model variance, as defined in Equation (4), is the difference between the average of the squared local drifts $\frac{1}{K}\sum\|\mathbf{u}_{t}^{(k)}\|_{2}^{2}$ and the squared global drift $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$.
+The first term reflects how far the individual worker models have moved–essentially, how much each worker has learned.
+The second term indicates how much of this learning is retained in the global model after aggregation.
+The interplay between these two quantities is crucial.
+For example, when the local drifts are high but the global drift is low, the variance increases, signaling the need for synchronization.
+This scenario suggests that while individual workers have made significant progress (as indicated by high local drifts), this progress is not being effectively captured in the global model (indicated by the low global drift).
+In other words, the worker models have moved significantly, but the global model has remained relatively stationary in this high-dimensional space.
+This misalignment indicates that training is no longer progressing optimally, as the workers are moving towards disparate and conflicting local minima, making it crucial to synchronize and realign them.
+Conversely, when both the local and global drifts are either low or high, synchronization is not necessary, and the variance naturally remains low.
+Neither the average of the local drifts nor the global drift alone provides a complete picture of the collective training progress.
+Relying solely on one or the other would lead to suboptimal synchronization decisions and likely prove ineffective.
+In FDA, it is the relationship between these quantities, as captured by the model variance, that offers valuable insights and guides the crucial decision of when to synchronize.
+SketchFDA vs.
+LinearFDA: Both methods send the squared norm of the drift $\|\mathbf{u}_{t}^{(k)}\|_{2}^{2}$, but differ in the additional accompanying lower-dimensional representation they transmit (Figure 2):
+  1. (1)
+    SketchFDA: An AMS sketch of the local drift.
+  2. (2)
+    LinearFDA: The dot product of a vector and the local drift.
+The key difference between these two variants lies in the fidelity of approximation of the model variance.
+While both methods conservatively overestimate the variance, SketchFDA provides a provably accurate estimation, which is expected to lead to fewer synchronizations.
+LinearFDA requires less computational effort and bandwidth to create and communicate the local states, but may overestimate variance by too much, causing unnecessary synchronizations.
+Figure 2. SketchFDA & LinearFDA: Local State structure.
+SketchFDA: Choice of $l$ and $m$.
+We empirically measured the approximation achieved with sketch dimensions of $l=5$ rows and $m=250$ columns (as defined in Section 3.1): these settings yield an error bound of $\epsilon\approx 6\%$ and a probabilistic confidence of $(1-\delta)\approx 95\%$.
+Based on our experiments, we have adopted these values in our experiments and recommend them.
+Using these values, the byte-size of a sketch is $l\cdot m\cdot 4\,\mathrm{bytes}=5\,\mathrm{kB}$, significantly smaller than the size of all our models.
+Sketches of smaller size could be used, albeit weakening the approximation of the variance.
+However, given that LinearFDA similarly weakens approximation and avoids using AMS sketches, in the interest of space we do not explore varying AMS sketch sizes in this paper.
+FDA: Asynchronous Operation.
+As mentioned in Section 2, FDA can be readily modified to operate asynchronously.
+In this setup, one worker-node acts as a coordinator, aggregating local states and determining whether synchronization is needed each time a local state is received.
+This decision is based on the most recent local states from all workers.
+It is important to note that, since local states are small in size, asynchronous operation is unlikely to alleviate bandwidth issues.
+The primary advantage is that it allows training to continue even in the presence of stragglers.
+Asynchronous operation might also be beneficial in rare cases where the overhead of initializing communication dominates the actual transmission time.
 4. Experiments
 4.1. Setup
+Table 2 provides a comprehensive overview of our experiments.
+For each experiment, we detail the Neural Network (NN) architecture, its parameter count ($d$), and the dataset used for training.
+The table also specifies key hyper-parameters: the batch size ($b$), the number of workers ($K$), and the FDA-specific variance threshold ($\Theta$).
+Additionally, we indicate the chosen optimizer (as detailed in Section 3) and the training algorithms employed for each configuration.
 Platform.
 We employ TensorFlow (Abadi et al., 2016), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments.
-We modified TensorFlow to include our FDA variants and all competitive algorithms.
+We used TensorFlow to implement our FDA variants and all competitive algorithms.
 All relevant code, figures, and data of this study are available in https://github.com/miketheologitis/FedL-Sync-FDA.
+Hardware & Infrastructure.
+We conducted our experiments on the ARIS High performance computing (HPC) environment [^2] https://www.hpc.grnet.gr/en/hardware-2/, utilizing a cluster of 44 GPU-accelerated worker-nodes.
+Each worker is equipped with two NVIDIA Tesla K40m GPUs and interconnected via an InfiniBand FDR14 network, providing up to 56 GB/s of bandwidth.
+Crucially, our evaluation remains agnostic to the underlying infrastructure of the specific workers.
 Table 2. Summary of Experiments
@@
 The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
-Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021).
+Lastly, we explore a transfer learning scenario on the dataset CIFAR-100 (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021).
 For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2021).
@@
 For example, comparing DDL algorithms based on the average cost of a training epoch can be misleading, as it does not consider the effects on the trained model’s quality.
-To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in the Figures).
+To achieve a comprehensive performance assessment of FDA, we define a training run as the process of executing the DDL algorithm under evaluation, on (a) a specific DL model and training dataset, and (b) until a final epoch in which the trained model achieves a specific testing accuracy (termed as Accuracy Target in figures).
 Based on this definition, we focus on two performance metrics:
@@
     Communication cost, which is the total data (in bytes) transmitted by all workers.
-Translating this cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers.
+Notably, communication cost is unaffected by the training data volume since only model updates (when synchronizing) and local states (at each step), but not training data, are transmitted.
+Thus, the communication cost mainly depends on the complexity (number of parameters) of the used model.
+Translating the communication cost to wall-clock time (i.e., the total time required for the computation and communication of the DDL) depends on the network infrastructure connecting the workers and on the overhead of establishing and initializing communication.
 Its impact is larger in FL scenarios, where workers often use slower Wi-Fi connections.
   2. (2)
-    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in the Figures) performed by each worker.
-Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the nodes.
-Its impact is larger for nodes with lower computational resources.
+    Computation cost, which is the number of mini-batch steps (termed as In-Parallel Learning Steps in figures) performed by each worker.
+Translating this cost to wall-clock time is determined by the mini-batch size and the computational resources of the worker-nodes.
+Its impact is larger for workers with lower computational resources.
 Hyper-Parameters & Optimizers.
@@
 LeNet-5 on MNIST.
-At Non-IID: Label ”0”, the samples of Label ”0” are assigned to few workers.
+At Non-IID: Label "0", the samples of Label "0" are assigned to few workers.
 At Non-IID: 60%, 60% of the dataset is sorted and allocated to workers, causing some workers to receive many samples from the same label
@@
 These KDE plots provide a high-level overview of the cost trade-off for training accurate models.
+The varying levels of opacity in the filled areas of the KDE plots represent the density of the underlying data points: higher opacity indicates areas with a greater concentration of data, whereas lower opacity signifies less dense areas.
 As an illustrative example, Figure 3 depicts the strategies’ bivariate distribution for the LeNet-5 model trained on MNIST with different data heterogeneity setups.
 In these plots, the SketchFDA distribution is generated from experiments across all hyper-parameter combinations ($\Theta$ and $K$ in Table 2) that attained the Accuracy Target of $0.985$.
+The observed high variance in the method’s distribution stems from the varying $K$ and $\Theta$ values.
+In subsequent subsections, we elucidate how these hyper-parameters influence the communication and computation costs.
 Figure 5. DenseNet121 on CIFAR-10
@@
 Figure 7.
-Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right).
+Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (top), and $0.78$ (bottom).
 Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively.
-A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model.
-Figure 7. LeNet-5 on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.98$
-Figure 8. VGG16* on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.994$
-Figure 9. DenseNet121 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.8$
-Figure 10. DenseNet201 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.78$
+A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
 FDA balances Communication vs.
@@
 Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost).
-This is evident in Figures 234, and 5 (where Synchronous appears in the bottom right — low computation, very high communication).
+This is evident in Figures 346, and 6 (where Synchronous appears in the bottom right — low computation, very high communication).
 Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation.
-Indeed, as shown in Figures 2-5, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation.
-Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 23,  4, and 5.
+Indeed, as shown in Figures 3-6, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation.
+Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 34,  6, and 6.
 In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency.
@@
 We first clearly notice this with VGG16* on MNIST in Figure 4 for all three data heterogeneity settings.
-For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication.
+For a 0.001 increase in accuracy (effectively 10 misclassified testing images), FedAdam needs approximately 2-7$\times$ more communication and 3-7$\times$ more computation, respectively.
+This can be seen by comparing the figures at the left column of Figure 4 with the corresponding ones in the right column.
 Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication.
 On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement.
-For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  4, and 5), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201).
+For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  6, and 6), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201).
 In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
@@
 Our empirical investigation reveals the FDA methods’ noteworthy resilience in such heterogeneous environments.
-For LeNet-5 on MNIST, as illustrated in Figure 2, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label ”00”, $60\%$).
-Similarly, for VGG16* on MNIST, Figure 3 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label ”00”, Label ”$8$”); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
-FDA generalizes better.
+For LeNet-5 on MNIST, as illustrated in Figure 3, the computation and communication costs required to attain a test accuracy of $0.985$ show negligible differences across the IID and the two Non-IID settings (Label "00", $60\%$).
+Similarly, for VGG16* on MNIST, Figure 4 demonstrates that achieving a test accuracy of $0.995$ incurs comparable computation and communication costs across the IID and the two Non-IID settings (Label "00", Label "$8$"); while overall costs are aligned, the distributions of the computation costs exhibit greater variability, yet remain closely consistent with the IID scenario.
+FDA has a lower generalization gap.
 The factors determining how well a DL algorithm performs are its ability to: (1) make the training accuracy high, and (2) make the gap between training and test accuracy small.
@@
 The patterns of performance we highlighted are consistent across most of the conducted tests.
+Figure 8. LeNet-5 on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.98$
+Figure 9. VGG16* on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.994$
+Figure 10. DenseNet121 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.8$
+Figure 11. DenseNet201 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.78$
 Dependence on $K$.
 In distributed computing, scaling up typically results in proportional speed improvements.
-In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup.
+In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup (Yu et al., 2019).
 This unpredictability is starkly illustrated with LeNet-5 and VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies.
-Figures 7, and 8 (left) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication.
+Figures 8, and 9 (top) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication.
 These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources.
-However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (left part of Figures 910), the expected behavior starts to emerge.
+However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (top part of Figures 1011), the expected behavior starts to emerge.
 Especially for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies.
@@
 Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers.
-Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 5-7.
+Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 8-11.
 Specifically, they require up to 30 times less communication than FedAdam, 4 times less than FedAvgM, and up to 2.5 orders of magnitude less than Synchronous.
@@
 A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence.
-This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 7-10).
-Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 9 and 10.
+This impact of $\Theta$ is consistently observed across both FDA strategies, and all learning tasks, and data heterogeneity settings (Figures 8-11).
+Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 10 and 11.
 It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
-Figure 11. Empirical Estimation of the Variance Threshold
-Figure 12. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
 FDA: Choice of $\Theta$.
-The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 7,  89, and 10.
+The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 8,  910, and 11.
 Therefore, identifying this range becomes crucial.
@@
 Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 2), thereby investigating different computation and communication trade-offs.
-For instance, in the ARIS-HPC environment (https://www.hpc.grnet.gr/en/hardware-2/) with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible.
+For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible.
 In such contexts, lower $\Theta$ values are favored due to their computational efficiency.
 On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
-To assist researchers in selecting the variance threshold, Figure 11 presents empirical estimations for $\Theta$ across three distinct learning settings: (1) FL (assuming a common channel of $0.5$Gbps), (2) Balanced communication-computation equilibrium setting, and (3) our HPC environment at the ARIS supercomputer.
+Figure 12. Empirical Estimation of the Variance Threshold
+To assist researchers in selecting the variance threshold, Figure 12 presents empirical estimations for $\Theta$ across three distinct learning settings:
+  1. (1)
+    FL, assuming a common channel of $0.5$Gbps, where
     $\Theta_{FL}$ $=4.91\cdot 10^{-5}\cdot d$
+  2. (2)
+    Balanced communication-computation equilibrium, where
     $\Theta_{B}$ $=3.89\cdot 10^{-5}\cdot d$
+  3. (3)
+    Our HPC environment at the ARIS supercomputer, where
     $\Theta_{HPC}$ $=2.74\cdot 10^{-5}\cdot d$
@@
 Sketch.
-In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 234, and 5.
+In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 346, and 6.
 This suggests that the precision of the variance approximation is not critical.
@@
 On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, if communication-efficiency is paramount.
-5. Conclusions
+Figure 13. ConvNeXtLarge on CIFAR-100 (transfer learning from ImageNet) — Deployment of FDA during the fine-tuning stage with Accuracy Target of $0.76$
+5. Conclusions and Future Work
 In this paper, we introduced Federated Dynamic Averaging (FDA), an innovative, adaptive and communication-efficient algorithm for distributed deep learning.
-Essentially, FDA makes informed, dynamic decisions on when to terminate training rounds based on approximations of the model variance.
+Essentially, FDA makes informed, dynamic decisions on when to synchronize the local models based on approximations of the model variance.
 Through extensive experiments across diverse datasets and learning tasks, we demonstrated that FDA significantly reduces communication overhead (often by orders of magnitude) without a corresponding increase in computation or compromise in model performance—contrary to the typical trade-offs encountered in the literature.
-Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates overfitting.
-Our results push the limits of modern communication-efficient distributed deep learning, paving the way for scalable, dynamic, and broadly applicable strategies.
+Furthermore, we showed that FDA is robust to data heterogeneity and inherently mitigates over-fitting.
+Our results push the limits of modern communication-efficient distributed deep learning, paving the way for more scalable, dynamic, and broadly applicable strategies.
+An interesting direction for future work is whether the value of $\Theta$ can be dynamically adjusted in order to achieve (or not to exceed) a target average bandwidth consumption.
+Since the expected behavior is that the communication cost decreases when $\Theta$ increases, such an approach seems feasible (i.e., increasing $\Theta$ when the bandwidth consumption is higher than what is desired), especially by using statistics.
+We plan to look into this extension in the future.
 Acknowledgements.
+We wish to thank the anonymous reviewers and the meta-reviewer for their insightful comments and suggestions.
 This work was supported by the EU project CREXDATA under Horizon Europe agreement No.